E12-5

A,B,C,D,E

On January 2, 2005cember 31, 2004, assum assuming to 3. EXERCISES, PROBLEMS, AND CASES TO ACCOMPANY FINANCIAL ACCOUNTING December 31, 2004? 200
Answer #1

a. Factors to be considered in determining whether the lease is a capital lease or an operating lease:

Capital lease is the lease that transfers substantially all the risks and rewards incident to ownership of an asset. In case of capital lease, the lessee records the leased asset as a fixed asset, depreciation expense for leased asset (over the lease term), liability for the present value of leased asset using discount rate i.e. incremental borrowing rate of lessee's.

In case of operating lease, lessee records each lease payment as operating expense.

b. Total expense recognized under operating lease $3000 every year.

c. If lease considered as capital lease, expense to be recognized over the three years:

Depreciation every year = Present value / lease term = 7460 / 3 = 2487 every year

Interest @ 10% = 7460 * 10% = 746

Difference = 9000 -7460 - 746 = 794 gain on account of lease

d. Finance lease will result in highest income in each of the three years since the transactions are recorded based on present value, depreciation and interest recorded. Total cash outflow $3000 every year.

e. Highest cash flow will be same for both types of lease as the lease payment is $ 3000 every year.